# Open-Source ESP32-C3 Wi-Fi Drivers for Static Analysis

### v2

@@ Abstract.
The Battery-Free Internet of Things will revolutionize our understanding of sustainable communication, as it operates on energy harvested from the environment (e.g., by solar cells). On these systems, harvested energy is dependent on unpredictable environmental conditions; therefore, device operations, including those of its networking stack, must be resilient to power failures. Reactive intermittent computing provides an approach for solving this by notifications of impending power failures, which is implemented by monitoring the harvested energy buffered in a capacitor. However, to use this power-failure notification and guarantee forward progress, systems must break down tasks into atomic sub-tasks (transactions) that can be predictably finished before the available energy runs out. Thus, static program-code analysis must determine the worst-case energy consumption (WCEC) of all transactions. In the case of Wi-Fi–capable devices, drivers are often closed-source, and it is therefore unclear whether their transactions have a WCEC since static analysis requires all code and its semantics.
In this work, we integrate a transactional networking stack with reverse-engineered Wi-Fi drivers to enable full-stack WCEC analysis for physical transmission and reception of packets. Further, we extended a static worst-case analysis approach with a resource-consumption model of our Wi-Fi driver. Our evaluation with the RISC-V–based ESP32-C3 platform gives worst-case bounds with our static analysis for the transactions of the full communications stack, therefore showing that Wi-Fi–based reactive intermittent computing is feasible.
@@ 1. Introduction
Battery-Free Systems
Battery-free devices (Ahmed et al., 2024) rely on energy harvested from various sources, such as light, radio-frequency signals, or vibrations. Being battery-free not only results in independence from battery requirements but also makes novel energy-self–sufficient applications possible: Battery-free medical implants (Kim et al., 2024) could reduce the need for surgical replacements due to degradation in battery performance, which puts such scenarios in the realm of both safety- and security-critical systems (Mottola et al., 2024). Naturally, the unpredictable nature of harvesting energy means that a started task can face a power failure before reaching a checkpoint, stored in non-volatile memory. These power failures cause two major problems: First, such a failure can leave the system in an inconsistent state. Second, when power failures occur during energy-intensive operations with transactional semantics (e.g., sending out a packet over a transceiver), the system might repeatedly face power failures, leaving the system in a livelock-like situation.
@@
To solve these two problems of inconsistency and starvation, the use of static worst-case program analysis can give guarantees on the execution of operations. Specifically, static worst-case energy consumption (WCEC) analysis techniques (Choi et al., 2022; Raffeck et al., 2024; Yarahmadi and Rohou, 2020) can determine a safe upper bound on the required energy demand of operations. These static program analysis techniques originally stem from the real-time systems community (Wilhelm et al., 2008) and their requirement to determine worst-case execution time (WCET) estimates for guaranteeing timeliness. With WCEC estimates and knowledge during runtime of the available energy in the storage (i.e., capacitor), operations run without the risk of power failures. Thus, consistency and forward progress is guaranteed.
@@
A necessary requirement for applying such static resource-consumption analyses is the availability of all code along with its resource-consumption–related semantics (Abella et al., 2015; Schoeberl, 2009). This required knowledge is particularly important for power-consuming operations, such as radio transceivers, since inaccuracies in the static behavior modeling can result in (1) unsound or (2) impractically pessimistic analysis estimates. Unfortunately, for our embedded target platform, the RISC-V–based ESP32-C3 system-on-chip (SoC) from Espressif, the Wi-Fi drivers are closed-source, and the Wi-Fi module’s behavior is undocumented on the granularity that we require for static analysis.
@@
Fortunately, sophisticated tooling support from the security community enabled us to reverse engineer the necessary knowledge we require for conducting worst-case analyses. Further, the existing ESP32-Open-MAC project (Devreker, 2023), targeting Tensilica’s Xtensa SoC ESP32, allows us to benefit from an existing infrastructure. In summary, this work’s contributions are threefold:
Case Study on Reverse Engineering: We present our development of a Wi-Fi stack for the ESP32-C3, relying on static reverse-engineering analysis techniques (i.e., Ghidra) combined with dynamic analysis (i.e., binary-patched executables). The stack is designed to be fully statically analyzable and therefore suited for intermittently-powered devices.
Energy-Consumption & Execution-Time Model for Worst-Case Analysis: Our static analysis comes with a fine-grained energy-consumption model of the Wi-Fi driver for the ESP32-C3 target platform.
Open-Source Implementation & Evaluation: Based on our publicly available driver for ESP32-C3, we present evaluation results that validate the stack’s property of being statically analyzable.
As part of this work, we envision that concerns from the security domain, such as security auditability of the Wi-Fi stack’s behavior, can be interconnected with questions from the static-analysis domain, such as the resource demand of the Wi-Fi stack’s behavior.
Paper’s Organization
Section 2 introduces the necessary background information for our domain of Wi-Fi–enabled battery-free systems. Our main problems are stated in Section 3, while Section 4 gives insight into how we solved these problems. Section 5 presents our evaluations on a widely used embedded platform (i.e., ESP32-C3). Related work is discussed in Section 6. Section 7 outlines future work and concludes the paper.
@@ 2. Background
The following section first gives some background regarding embedded network stacks (see Section 2.1). In section 2.2, we discuss mechanisms through which processors and Wi-Fi peripherals communicate in low-power SoCs. Finally, 2.3 describes the approach of worst-case resource-consumption analysis and its role in intermittent systems.
@@ 2.1. Network stack
The network stack on embedded devices typically consists of a lightweight TCP/IP implementation and a driver that interacts with the physical network interface (e.g., Wi-Fi peripheral). At the highest level of the stack is the application that provides data to be transmitted, while at the lowest level is the physical layer (PHY), which sends out the data onto the physical medium. On top of the physical layer is the medium access control (MAC) that coordinates transmission between multiple devices. The MAC layer is tasked with network association, collision avoidance, or data encapsulation. In the case of Wi-Fi, collision avoidance (CSMA/CA) mechanism is used to prevent multiple devices from transmitting at the same time. Each layer of the network stack encapsulates data from the upper layer based on network configuration and passes it to the layer below. On intermittent devices, power may be lost at any point before packet transmission is completed, leading to loss of progress up to that point.
In the case of Wi-Fi, PHY operations are implemented in hardware due to strict timing requirements. Wi-Fi peripherals that use the FullMAC architecture may implement both MAC and PHY layer operations in hardware. In some peripherals, tasks such as network association may be implemented in software (SoftMAC).
Popularly, Wi-Fi devices operate in the basic service set (BSS) topology and connect to an access point (AP) that arbitrates communication within the network. Once connected, the devices are known as stations.
@@ 2.2. Wi-Fi Hardware
High-speed communication between the Wi-Fi peripheral and the processor within an SoC is typically carried out using a combination of memory-mapped I/O (MMIO) and direct memory access (DMA) (Schulz, 2018). The Wi-Fi peripheral is configured through MMIO, while data transfer for transmission and reception is performed using DMA. DMA allows the CPU to remain free for other tasks instead of handling the transfer. Espressif devices such as ESP32-C3 follow this architecture. The peripheral communicates events such as successful transmission or reception of a packet through interrupts. On the other hand, external Wi-Fi chipsets may use the SPI interface for communication with the host controller. Besides the ESP32-C3’s Wi-Fi chip, the CYW43439 is another example suitable for low-power systems (Cypress Semiconductor Corporation, 2021).
2.3. Static Worst-Case Analysis
Static worst-case resource-consumption analysis comprises two major pillars: (1) a hardware-agnostic path analysis and (2) a hardware-dependent cost modeling capturing the actual resource demand (i.e., execution time or energy demand) on the target platform. The main idea is to formulate a mathematical problem formulation, usually an integer linear program (ILP): The ILP’s constraints originate from the path analysis. The ILP’s objective function states the maximization of the actual costs with the given path constraints. Eventually, the ILP’s solution yields a mathematically proven worst-case bound, which, in turn, is necessary to ensure runtime guarantees.
@@ 3.1. Problem of Closed-Source Wi-Fi Drivers
Our approach to intermittently-powered devices exploits static worst-case analysis techniques to gain knowledge on the worst-case energy consumption of operations. During runtime, our system’s scheduler checks the available energy and dispatches tasks only if sufficient energy is available to guarantee freedom from power failures. This WCEC-aware approach gives both memory-consistency and forward-progress guarantees. The trade-off we exploit is favoring expensive design-time analyses over complicated runtime checkpointing mechanisms. Once deployed, systems achieve provable resilience against power failures. However, these design-time analyses require the code to be available. Our target platform is the ESP32-C3, a RISC-V–based, widely used, and low-cost SoC. Although BLE, likewise available on the ESP32-C3, is less power-intensive compared to Wi-Fi, we are interested in the more challenging Wi-Fi stack being of high relevance: In 2024, an estimated 31 % of IoT devices used Wi-Fi for connectivity (Sinha, 2024). For this ESP32-C3 platform, the Wi-Fi drivers were not available as analyzable source code, which is a common problem for embedded SoC platforms. Of particular relevance for our analysis are the MAC and PHY layer operations.
Our Approach: To solve the problem of the closed source, our approach exploits several reverse-engineering techniques on Espressif’s closed-source Wi-Fi driver. Fortunately, Espressif legally allows reverse engineering this driver. Based on the reverse-engineering results, we contribute an open-source driver for the ESP32-C3.
@@ 3.2. Missing Resource-Consumption Model
Although the developed open-source driver gives us the freedom to perform static analysis,
a major problem is connecting memory-mapped IO operations to semantics with regard to the power demand.
When considering the following code for modifying memory
*((uint32_t*)0x60033ca0) &= 0xff00efff
the static analysis is not aware of the semantics of this memory access.
On the ESP32-C3 target platform, this operation switches on the transceiver from a sleep state to standby mode (see also Figure 1).
As a consequence thereof, our static analysis of this memory access requires the notion that the power drawn increases after this operation.
Similarly, specific instructions also require a longer execution time (than a normal RAM access), which in turn leads to higher energy demand (i.e., power over time).
Consequently, we require a model that comprehensively accounts for power, energy, and time.
Our Approach: For our target, we present a resource-consumption model that considers transitions in the hardware that change the power demand of the entire SoC. Our analysis makes use of this model and tracks program paths in a context-sensitive way.
@@ 4. Approach
We use a combination of commonly available static and dynamic reverse-engineering analysis methods for identifying the registers and correspondingly their contents. Espressif provides their SoftMAC implementation as binaries under the permissive Apache license. The binaries also contain the symbol table.
@@ 4.1. Reverse Engineering
Through decompilation, the execution flow and accordingly the registers used for transmission and reception can be identified. Functions within the binaries are named based on their purpose, which leads to informative decompilation results. Additionally, the interrupt configuration used by the Wi-Fi peripheral can be identified as this is not exposed to developers.
However, decompilation alone is insufficient to identify control and data flows. Several of the registers are configured based on data that is internal to the driver, whose structure is not documented. Moreover, the peripheral also has several slots for transmission, which is also dynamically selected based on the network configuration. Therefore, we use QEMU to dynamically determine the register contents as well as the slots.
@@
The QEMU version in (lcgamboa, 2023), a fork of Espressif’s QEMU, supports Wi-Fi for ESP32-C3 by identifying a select set of registers and emulating an open Wi-Fi AP. We modify this fork to log all peripheral accesses and their contents. This allows us to verify register addresses as well as determine the register operations corresponding to the Wi-Fi network configuration in QEMU.
While QEMU provides insights into all the peripheral’s register accesses and their contents, its emulation does not fully account for peripheral configuration. It only emulates the DMA transfer as well as the interrupt mechanism. Moreover, it is not always possible to deduce the purpose of the register operation.
Binary Patching for Injecting Tracing Code
We modify the control flow to trace execution in real-time. We make no modifications to the binaries themselves but replace exported symbol names within some binary files. The new symbol name can then be declared as a function as part of a regular ESP-IDF sketch and can call the original function with the provided arguments. This also allows us to peek into the arguments themselves without using breakpoints or external debuggers. The resulting binary can also be run on the emulator or on the ESP32-C3. However, this only allows us to create such function redirects only for those functions that are declared in one object file but called from another.
These reverse-engineering methods in combination with the existing infrastructure in ESP32-Open-MAC yield necessary information for constructing the Wi-Fi driver for ESP32-C3.
[device graph]device graph showing energy states
@@ 4.2. The Open-Source ESP32-C3 Driver
Once the registers and transmit/receive mechanisms are identified, we use the software architecture shown in Figure 2 to develop the driver.
@@
The driver task interacts with the MMIO registers. It also manages the memory allocated to DMA structures required by the peripheral. All memory is statically allocated to aid WCET/C analysis. Processing of interrupts is deferred by the interrupt handler to the driver task. Incoming 802.11 packets are passed to the MAC layer for further processing. The driver task also defines two queues: a hardware event queue and a MAC event queue. The interrupt handler notifies the driver of the occurrence and reason for the interrupt using the hardware event queue. The MAC task also writes packets to be transmitted into this queue. On the other hand, the driver task writes any received packets into the MAC event queue to signal the MAC task of received packets.
@@ 4.3. Building a Resource-Consumption Model
For tasks such as transmission and reception, the Wi-Fi peripheral functions independently of the CPU. Therefore, static code analysis has to be supplemented with resource consumption models, namely time and energy.
@@
The Wi-Fi peripheral performs packet transmission and notifies the CPU of the status through interrupts. Similarly, some MMIO register operations take time due to ongoing receive tasks. However, these tasks are still completed or fail within a finite amount of time and therefore have upper bounds on their execution times.
@@
The device energy model should consist of all possible energy states, as well as the corresponding energy consumption. In our current setup, we assume that the CPU is in the always-on state; however, modeling the ESP32-C3’s sleep states, (low-power) run states, and frequency changes is possible (Dengler et al., 2023). For the Wi-Fi peripheral, we define three states of operation: Sleep, Standby and Transmitting. The Wi-Fi peripheral is powered-down in and therefore consumes no energy. In , the Wi-Fi peripheral is powered up and may receive packets, thereby consuming energy. In , the device transmits packets. The energy consumption varies based on the duration of transmission. No reception occurs in this state, as the peripheral owns the channel through the CSMA-CA mechanism. The device state machine is shown in Figure 1.

### v3

# Reverse Engineering the ESP32-C3 Wi-Fi Drivers for Static Worst-Case Analysis of Intermittently-Powered Systems

@@ Abstract.
The Internet of Batteryless Things revolutionizes sustainable communication as it operates on harvested energy. This harvested energy is dependent on unpredictable environmental conditions; therefore, device operations, including those of its networking stack, must be resilient to power failures. Reactive intermittent computing provides an approach for solving this by notifications of impending power failures, which is implemented by monitoring the harvested energy buffered in a capacitor. However, to use this power-failure notification and guarantee forward progress, systems must break down tasks into atomic transactions that can be predictably finished before the energy runs out. Thus, static program-code analysis must determine the worst-case energy consumption (WCEC) of all transactions. In Wi-Fi–capable devices, drivers are often closed-source, which avoids the determination of WCEC bounds for transactions since static analysis requires all code along with its semantics.
In this work, we integrate an energy-aware networking stack with reverse-engineered Wi-Fi drivers to enable full-stack WCEC analysis for physical transmission and reception of packets. Further, we extended a static worst-case analysis tool with a resource-consumption model of our Wi-Fi driver. Our evaluation with the RISC-V–based ESP32-C3 platform gives worst-case bounds with our static analysis approach for the transactions of the full communication stack, therefore showing that Wi-Fi–based reactive intermittent computing is feasible.
@@ 1. Introduction
Batteryless Systems
Batteryless devices (Ahmed et al., 2024) rely on harvested energy. Being batteryless not only results in independence from battery requirements but also makes novel energy-self–sufficient applications possible: Batteryless medical implants (Kim et al., 2024) reduce the need for surgical replacements due to degradation in battery performance, putting them into the scope of safety- and security-critical systems (Mottola et al., 2024). Naturally, the unpredictable nature of harvesting energy means that a started task can face a power failure before reaching a checkpoint stored in non-volatile memory. Power failures cause two major problems: First, failures can leave the system in an inconsistent state. Second, with failures during energy-intensive operations with transactional semantics (e.g., sending packets), systems might repeatedly face failures without making progress.
@@
To address inconsistency and starvation, static worst-case program analysis can guarantee power-failure–free execution in intermittent systems. Specifically, static worst-case energy consumption (WCEC) analysis (Choi et al., 2022; Raffeck et al., 2024; Wägemann et al., 2018; Yarahmadi and Rohou, 2020) determines a safe upper bound on operations’ energy demand. These analysis techniques originally stem from determining the worst-case execution time (WCET) to guarantee timeliness in real-time systems (Wilhelm et al., 2008). With WCEC estimates and the available energy in the storage (i.e., capacitor) during runtime, operations execute without the risk of power failures. Consequently, consistency and forward progress are guaranteed under consideration of the currently available energy.
@@
A necessary requirement for such static resource-consumption analyses is the availability of all code along with its resource-consumption semantics (Abella et al., 2015; Schoeberl, 2009). This knowledge is particularly important for power-consuming operations, such as radio transceivers, since inaccuracies in the static behavior modeling can result in unsound or impractically pessimistic analysis estimates. Unfortunately, for our target platform, the RISC-V–based ESP32-C3 system-on-chip (SoC), the Wi-Fi drivers are closed-source, and the Wi-Fi peripheral’s behavior is undocumented.
@@
The ESP32-Open-MAC project (Devreker, 2023) provides reverse-engineered Wi-Fi drivers for an Xtensa-based SoC, which we leveraged for the RISC-V–based ESP32-C3. Our work presents open-source Wi-Fi drivers for the ESP32-C3 alongside their static worst-case analyzability. In summary, our contributions are threefold:
Case Study on Reverse Engineering: We present our reverse engineering efforts for the Wi-Fi driver of the ESP32-C3.
Energy-Consumption & Execution-Time Model for Worst-Case Analysis: Our static analysis comes with a fine-grained energy-consumption model.
Open-Source Implementation & Evaluation: Using the reverse-engineered knowledge, we develop an open-source Wi-Fi driver for the ESP32-C3 and present evaluation results that validate the driver’s property of being statically analyzable.
@@ 2. Background
This section gives background on network stacks and mechanisms for processor and Wi-Fi module communication. Finally, we outline worst-case resource-consumption analyses of program code.
@@ 2.1. Network Stack
The network stack on embedded devices typically consists of a lightweight TCP/IP implementation and a driver that interacts with the physical network interface (e.g., Wi-Fi peripheral). At the highest level of the stack is the application that provides data to be transmitted, while at the lowest level is the physical layer (PHY), which sends out the data onto the physical medium. On top of the physical layer is the medium access control (MAC) that coordinates transmission between multiple devices. In the case of Wi-Fi, the carrier sense multiple access/collision avoidance (CSMA/CA) mechanism is used to coordinate transmissions. Each layer of the network stack encapsulates data from the upper layer and passes it to the layer below. On intermittent devices, power may be lost at any point before packet transmission is completed, leading to a loss of progress up to that point.
In the case of Wi-Fi, PHY-layer operations are implemented in hardware due to strict timing requirements. Some Wi-Fi peripherals implement MAC operations in software (SoftMAC), while others implement both MAC and PHY operations in hardware (FullMAC).
@@ 2.2. Wi-Fi Hardware
High-speed communication between the Wi-Fi peripheral and the processor within an SoC is typically carried out using a combination of memory-mapped I/O (MMIO) and direct memory access (DMA) (Schulz, 2018). The Wi-Fi peripheral is configured through MMIO, while data transfer for transmission and reception is performed using DMA. DMA allows the CPU to remain free for other tasks instead of handling the transfer. Espressif devices such as ESP32-C3 follow this architecture. The peripheral communicates events such as the successful transmission or reception of a packet through interrupts. On other platforms, external Wi-Fi chipsets may use the SPI interface to communicate with the host controller.
2.3. Static Worst-Case Resource Analysis
Static worst-case resource-consumption analysis of program code comprises two major pillars: (1) a hardware-agnostic program path analysis and (2) a hardware-dependent cost modeling capturing the actual resource demand (i.e., execution time or energy demand) for the actual target platform. The main idea is to formulate a mathematical problem formulation, usually an integer linear program (ILP): The ILP’s constraints originate from the path analysis. Abstract-interpretation techniques are further useful for refining the path constraints in a context-sensitive manner. The ILP’s objective function states the maximization of the actual costs with the given path constraints. Eventually, the ILP’s solution yields a mathematically proven worst-case bound. With regard to energy consumption, the WCEC bound is useful to guarantee the execution without power failures during the runtime of batteryless devices.
@@ 3.1. Problem of Closed-Source Wi-Fi Drivers
In intermittently-powered devices, static worst-case analysis techniques can be crucial to gain knowledge on the worst-case energy consumption of operations (Choi et al., 2022; Raffeck et al., 2024; Yarahmadi and Rohou, 2020). During runtime, the system’s scheduler checks the available energy (comparable to just-in-time checkpointing) and reactively dispatches tasks only if sufficient energy is available to guarantee freedom from power failures. The WCEC-aware approaches give both memory-consistency and forward-progress guarantees. The trade-off we exploit favors expensive design-time analyses over complicated runtime checkpointing mechanisms. Once deployed, systems achieve provable resilience against power failures. However, these design-time analyses require the code to be available. Our target platform, the ESP32-C3, is a RISC-V–based, widely used SoC from Espressif, which has closed-source Wi-Fi drivers. Although less power-intensive networking technology exists (e.g., BLE), we are interested in the Wi-Fi stack being of high relevance: In 2024, an estimated 31 % of IoT devices used Wi-Fi for connectivity (Sinha, 2024). Further, the recent article on The Internet of Batteryless Things (Ahmed et al., 2024) calls for action and states that popular network protocols, including Wi-Fi, should be adapted for intermittent operation. Unfortunately, for our target, the generic Wi-Fi drivers were not available as analyzable source code, which is a common problem in embedded SoCs. Of main interest for our analysis are the MAC and PHY layers.
Our Approach in a Nutshell: To solve the problem of the closed source, we exploit several reverse-engineering techniques on Espressif’s closed-source Wi-Fi driver alongside utilizing existing knowledge (Devreker, 2023). Fortunately, Espressif legally allows reverse engineering this driver. Based on the reverse-engineering results, we contribute an open-source driver for the ESP32-C3.
@@ 3.2. Missing Resource-Consumption Model
Although the developed open-source driver gives us the freedom to perform static analysis, a major problem is connecting memory-mapped IO operations to semantics with regard to the power demand.
When considering the following MMIO operation:
*((uint32_t*)0x60033ca0) &= 0xff00efff,
the static analysis is not aware of its semantics. On the ESP32-C3, this operation switches the transceiver from a sleep state to standby mode (see also Figure 2).
Hence, our static resource analysis needs to include the notion that the power drawn increases after this operation.
Similarly, specific instructions also require a longer execution time than a normal RAM access, which in turn leads to higher energy demand (i.e., power over time).
Consequently, a comprehensive model needs to account for power, energy, and time.
Our Approach in a Nutshell: For our target, we present a resource-consumption model that considers transitions in the hardware, which change the SoC’s power demand. Our analysis makes use of this model and tracks program paths in a context-sensitive way.
@@ 4. Approach
We use a combination of commonly available static and dynamic reverse-engineering analysis methods for identifying the registers and, correspondingly, their contents.
@@ 4.1. Reverse Engineering
Through decompilation, the program flow and, accordingly, the registers used for transmission and reception can be identified. Functions within the binaries are named based on their purpose, which leads to informative decompilation results. Additionally, the undocumented interrupt configuration used by the Wi-Fi peripheral can be identified. However, decompilation alone is insufficient to identify control and data flows. Several of the registers are configured based on data that is internal to the driver, whose structure is not documented. Moreover, the peripheral also has several slots for transmission, which are dynamically selected based on the network configuration. Therefore, we use QEMU to dynamically determine the register contents as well as the slots.
@@
The QEMU version in (lcgamboa, 2023), a fork of Espressif’s QEMU, supports Wi-Fi for ESP32-C3 by emulating an open Wi-Fi access point (AP). In particular, it only implements registers required for triggering a transmission and reading received packets through DMA as well as the status register, which encodes the reasons for an interrupt occurrence. We further modify this fork to log all register accesses and their contents within the Wi-Fi peripheral’s address range. While the QEMU fork provides insights into accesses of the peripheral’s registers and their contents used during the emulation, it does not fully emulate the hardware. Hardware initialization, for example, is not emulated. Moreover, it is not always possible to deduce the purpose of the register operation.
Binary Hooking for Injecting Tracing Code
The modules of Espressif’s Wi-Fi driver are provided as archives (.a) of object (.o) files. These object files have undefined symbols referring to functions defined in other modules, such as the binary blobs of the ESP32-C3 ROMs. Their definition is resolved during the final linking of the executable. To trace the execution, we hook into these symbols by replacing the symbol names and exploiting the original symbol names in our custom firmware to identify the control flow.
These reverse-engineering methods, in combination with the existing infrastructure in ESP32-Open-MAC, yield the necessary information for constructing the Wi-Fi driver for the ESP32-C3.
@@ 4.2. The Open-Source ESP32-C3 Driver
Once the registers and transmit/receive mechanisms are identified, we use the architecture shown in Figure 1 to develop the driver.
@@
The driver task interacts with the MMIO registers and manages memory allocated to DMA structures required by the peripheral. All memory is statically allocated to aid WCET/WCEC analysis. Processing of interrupts is deferred by the interrupt handler to the driver task. Incoming 802.11 packets are passed to the MAC task for further processing through a queue. The interrupt handler notifies the driver of the occurrence and reason for the interrupt using a hardware event queue. The MAC task also writes packets to be transmitted into this queue. On the other hand, the driver task writes any received packets into the MAC event queue to signal the MAC task of received packets.
@@ 4.3. Building a Resource-Consumption Model
For device operations such as transmission and reception, the Wi-Fi peripheral operates independently of the CPU. To yield sound results, static resource analysis requires a cost model specifying the maximum time and energy demands of operations. These paragraphs provide a generic overview of the cost model, while Section 5.2 discusses concrete values.
@@
The necessary timing-related costs comprise both protocol- and hardware-specific aspects. The application-specific protocol influences the maximum duration of device operations (e.g., through the maximum packet size). Additionally, timeouts and operation windows in the protocol allow deriving an upper bound on operations’ execution times, such as waiting to receive a packet. Bounds on the duration of operations also enable bounding the occurrence of interrupts related to the completion of a transmission or an incoming packet. The second aspect, the hardware-specific behavior, concerns the Wi-Fi peripheral’s MMIO interface. The execution time of certain MMIO operations, such as waiting for a bit to change its value, which signals the completion of a configuration update, depends on the reaction of the hardware.
[device graph]device graph showing energy states
@@
The energy-related cost model for the Wi-Fi peripheral consists of all possible device states with distinct power consumption. A simplified version of the device graph modeling the states and their transitions for the Wi-Fi peripheral is shown in Figure 2. The depicted graph features three states of operation: Sleep, Standby, and Transmitting. The Wi-Fi peripheral is powered down and inoperative in the sleep state ( ). In the standby state ( ), the Wi-Fi peripheral is powered up and may receive packets. Lastly, in the transmitting state ( ), the device actively transmits a packet. No reception occurs in this state, as the peripheral owns the channel through the CSMA/CA mechanism. For this paper, we assume that the CPU is always on and uses only one run mode, although modeling the ESP32-C3’s sleep states, run modes, and clock-frequency changes is generally possible (Dengler et al., 2023).